OpenStudy (anonymous):

whats the slope of a line perpendicular to the line that passes through (3,9) and (7,15)

OpenStudy (aroub):

15-9/7-3 6/4 3/2 y=mx+b , m=-2/3 , (3.9) 9=-2/3(3)+b 9=-2+b 11=b y=-2/3x+11